OpenStudy (anonymous):

help please algebra 2... 7n-4(5-2n)=10

OpenStudy (destinymasha):

Simplifying 7n + -4(5 + -2n) = 10 7n + (5 * -4 + -2n * -4) = 10 7n + (-20 + 8n) = 10 Reorder the terms: -20 + 7n + 8n = 10 Combine like terms: 7n + 8n = 15n -20 + 15n = 10 Solving -20 + 15n = 10 Solving for variable 'n'. Move all terms containing n to the left, all other terms to the right. Add '20' to each side of the equation. -20 + 20 + 15n = 10 + 20 Combine like terms: -20 + 20 = 0 0 + 15n = 10 + 20 15n = 10 + 20 Combine like terms: 10 + 20 = 30 15n = 30 Divide each side by '15'. n = 2 Simplifying n = 2

OpenStudy (anonymous):

n=2

OpenStudy (destinymasha):

maybe it'll help you

OpenStudy (anonymous):

The sides of a rectangle are in a 3:4 ratio. What is the length of the shorter sides of the rectangle if its perimeter is 42 cm?